Let X₁,X₂ ,...,Xₙ be a random sample from a Binomial distribution with parameters n and p, both unkown. Obtain estimators of n and p, using method of moments.
1
Expert's answer
2019-04-01T08:27:25-0400

Since we know that


"\text{E}[x]=np,"

"\text{Var}[x]=np(1-p),"

we can express "p" and "n" so that they are


"p=1-\frac{\text{Var}[x]}{\text{E}[x]},""n=\frac{\text{E}[x]}{p}=\frac{\text{E}^2[x]}{\text{E}[x]-\text{Var}[x]}."

On the other hand, we know that sample mean and sample variance are:


"\text{E}[x]=x_\mu=\frac{1}{n}\sum_{i=1}^{n}x_i,"


"\text{Var}[x]=S(x)=\frac{1}{n}\sum^{n}_{i=1}(x_i-x_\mu)^2,"

hence


"p=1-\frac{\frac{1}{n}\sum^{n}_{i=1}(x_i-x_\mu)^2}{\frac{1}{n}\sum_{i=1}^{n}x_i},"

"n=\frac{(\frac{1}{n}\sum_{i=1}^{n}x_i)^2}{\frac{1}{n}\sum_{i=1}^{n}x_i-\frac{1}{n}\sum^{n}_{i=1}(x_i-x_\mu)^2}."